# Mind-to-Image: Projecting Visual Mental Imagination of the Brain from fMRI

### v4

@@ 5.3 Can we reconstruct images from strong imagination / pure imagination?
Figure 4: Results from strong imagination. The subject is given an instruction on the fMRI screen, such as ”Imagine a portrait representing optimism”. Then, the subject purely imagines such image, giving an oral description. Then, the associated brain data is fed to the fMRI-to-Image model to produce a reconstruction, which is then evaluated compared to the oral description.
@@
Determining the content accuracy of images generated from strong imagination poses a unique challenge, as the ’ground truth’ images solely exist in the subject’s mental imagery, inaccessible for direct comparison. To navigate this, we based our evaluation on subjects’ descriptions of their imagined scenes, collected during or following the scanning sessions. The variability in the reconstructed content ranges from highly accurate renditions of specific imagined elements, like a tree, dark-haired woman, or flowers, to more abstract interpretations or plain errors. This also reflects the subjective nature of this task. Given the reliance on subjective descriptions and the inherent open-endedness of strong imagination, the results should be interpreted with caution. While it’s challenging to draw definitive conclusions about the model’s capacity for precise content reconstruction in this context, the instances of recognisable elements emerging in the reconstructions indicate that what could be achievable with increased data volume and further model refinement.
@@ 6.1 Ethical considerations
As we explore further into the capabilities of decoding and reconstructing visual content from brain activity, we get closer to the boundaries of personal thought and imagination, a concept known as ”mind privacy”. The ability to visualise an individual’s mental imagery raises questions about consent, mental autonomy, and the potential for misuse of such technologies. On the one hand, such technology provides potential for good, opening avenues for novel communication methods for those unable to speak or write, enhancing creative expression, and deepening our understanding of the human brain and its processes. On the other hand, it introduces the possibility of invasive breaches of personal thought, where the most private mental images and ideas could be exposed without consent. This requires a careful, deliberate approach to ensure that the development and application of these technologies are governed by ethical principles that prioritise individual privacy and autonomy. Currently, the barriers to entry, including the need for explicit consent and the use of expensive and sophisticated MRI machinery, provide a layer of protection.
@@ 6.2 Future work
Future directions for this line of research include the acquisition of more extensive data to improve the accuracy and specificity of content match in the reconstructed images. A refined experimental protocol for both weak and strong imagination tasks could lead to insights into the brain’s imaging processes, while improved evaluation methods for strong imagination would enhance our understanding of the model’s effectiveness in capturing the depth and detail of imagined content. Additionally, the question of whether MRI technology can be replaced or complemented by EEG is promising. Recent studies suggest the feasibility of reconstructing visual stimuli from brain activity using EEG (Benchetrit et al., 2023), which offers a more accessible, and cost-effective alternative to MRI.
@@ 7 Conclusion
Acknowledgements
The author acknowledges the support of the French Agence Nationale de la Recherche (ANR) under reference ANR-23-CE23-0023.
@@ Appendix A Uncurated samples for weak imagination
Figure 5: Uncurated results for weak imagination on the validation dataset. In each example, we present an image from the validation set of our surrealism image dataset, which is recollected from memory by the subject in the fMRI scanner, and the reconstructed image from the fMRI-to-Image model based on the brain data associated to the recollection of the original image.

### v5

@@ 5.3 Can we reconstruct images from strong imagination / pure imagination?
Determining the content accuracy of images generated from strong imagination poses a unique challenge, as the ’ground truth’ images solely exist in the subject’s mental imagery, inaccessible for direct comparison. To navigate this, we based our evaluation on subjects’ descriptions of their imagined scenes, collected during or following the scanning sessions. The variability in the reconstructed content ranges from highly accurate renditions of specific imagined elements, like a tree, dark-haired woman, or flowers, to more abstract interpretations or plain errors. This also reflects the subjective nature of this task. Given the reliance on subjective descriptions and the inherent open-endedness of strong imagination, the results should be interpreted with caution. While it’s challenging to draw definitive conclusions about the model’s capacity for precise content reconstruction in this context, the instances of recognisable elements emerging in the reconstructions indicate what could be achievable with increased data volume and further model refinement.
@@ 6.1 Ethical considerations
As we explore further into the capabilities of decoding and reconstructing visual content from brain activity, we get closer to the boundaries of personal thought and imagination, a concept known as "mind privacy". The ability to visualise an individual’s mental imagery raises questions about consent, mental autonomy, and the potential for misuse of such technologies. On the one hand, such technology provides potential for good, opening avenues for novel communication methods for those unable to speak or write, enhancing creative expression, and deepening our understanding of the human brain and its processes. On the other hand, it introduces the possibility of invasive breaches of personal thought, where the most private mental images and ideas could be exposed without consent. This requires a careful, deliberate approach to ensure that the development and application of these technologies are governed by ethical principles that prioritise individual privacy and autonomy. Currently, the barriers to entry, including the need for explicit consent and the use of expensive and sophisticated MRI machinery, provide a layer of protection.
@@ 6.2 Future work
Future directions for this line of research include the acquisition of more extensive data to improve the accuracy and specificity of content match in the reconstructed images. A refined experimental protocol for both weak and strong imagination tasks could lead to insights into the brain’s imaging processes, while improved evaluation methods for strong imagination would enhance our understanding of the model’s effectiveness in capturing the depth and detail of imagined content.
Additionally, the question of whether MRI technology can be replaced or complemented by EEG is promising. Recent studies suggest the feasibility of reconstructing visual stimuli from brain activity using EEG [3], which offers a more accessible, and cost-effective alternative to MRI.
Finally, in our experiments, the subject clearly expressed that the imagined images were getting clearer and more precise as he repeatedly attempted to imagine the same image. This opens a potentially promising avenue for "training" one’s own mental imagery, which could be applied to subjects presenting various degrees of aphantasia [24], i.e. the degree of their inability to create mental imagery. A set of experiments directed toward this question is thus a potential future work.
@@ Appendix A Uncurated samples for weak imagination
Fig. 5: Uncurated results for weak imagination on the validation dataset. In each example, we present an image from the validation set of our surrealism image dataset, which is recollected from memory by the subject in the fMRI scanner, and the reconstructed image from the fMRI-to-Image model based on the brain data associated to the recollection of the original image.
Appendix B Experimental details
All our models are trained on a single A100 GPU for 500 epochs with a batch size of 16, following [16] and our constraint of having bigger MLPs in the BrainNetwork class ($h$ was reduced to 2048). Trainings took anywhere from 15 to 25 hours depending on the size of the mask. The optimizers and training routine remained unchanged.